# Context-Picker: Dynamic context selection using multi-stage reinforcement learning

### v1

@@ Abstract
In long-context question answering (LCQA), determining the optimal amount of context for a given query is a significant challenge. Including too few passages may omit critical information, while including too many can introduce noise and reduce the quality of the answer. Traditional approaches, such as fixed Top-$K$ retrieval and single-stage reranking, face the dilemma of selecting the right number of passages. This problem is particularly pronounced for factoid questions, which often require only a few specific pieces of evidence. To address this issue, we introduce Context-Picker, a reasoning-aware framework that shifts the paradigm from similarity-based ranking to minimal sufficient subset selection. Context-Picker treats context selection as a decision-making process optimized via a human-inspired, two-stage reinforcement learning schedule: a recall-oriented stage that prioritizes the coverage of reasoning chains, followed by a precision-oriented stage that aggressively prunes redundancy to distill a compact evidence set. To resolve reward sparsity, we propose an offline evidence distillation pipeline that mines "minimal sufficient sets" via a Leave-One-Out (LOO) procedure, providing dense, task-aligned supervision. Experiments on five long-context and multi-hop QA benchmarks demonstrate that Context-Picker significantly outperforms strong RAG baselines, achieving superior answer accuracy with comparable or reduced context lengths. Ablation studies indicate that the coarse-to-fine optimization schedule, the redundancy-aware reward shaping, and the rationale-guided format all contribute substantially to these gains.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a standard paradigm for extending Large Language Models (LLMs) beyond their parametric knowledge, especially on knowledge-intensive and long-context question answering (LCQA) tasks (lewis2021retrievalaugmentedgenerationknowledgeintensivenlp; guu2020realmretrievalaugmentedlanguagemodel; izacard2022atlasfewshotlearningretrieval). By retrieving passages from an external corpus and conditioning generation on them, RAG mitigates hallucination and enables access to up-to-date or domain-specific information. In practice, most systems adopt a simple fixed Top-$K$ strategy: a retriever ranks candidate passages and the top-$K$ are concatenated and fed to the generator. However, the core design question of how much external context should be retrieved for a given query remains largely underexplored. When $K$ is too small, the model may miss critical evidence and break multi-hop reasoning chains, while an overly large $K$ introduces many weakly related passages, increasing inference cost and degrading answer quality through distractors, attention dilution, and the “lost-in-the-middle” phenomenon where LLMs under-utilize information placed in the middle of long prompts (liu2023lostmiddlelanguagemodels). Moreover, our experiments in Figure 1 show that increasing retrieval depth monotonically improves recall but leaves accuracy almost unchanged, which consistent with recent observations on long-context limitations in RAG (jin2024longcontextllmsmeetrag). This suggests that context handling in LCQA should be viewed not purely as a ranking problem, but as a subset selection problem: for each query, the system should construct a compact, query-specific evidence set that is sufficient for answering the question, rather than a long prefix of a ranked list.
@@
Figure 1: Accuracy vs. retrieval depth (Top-$K$) in a standard RAG pipeline. Recall increases with $K$, but answer accuracy does not improve, which is also reported in recent long context studies (jin2024longcontextllmsmeetrag).
@@
Recent works on retrieval-augmented generation solve this problem from two main directions. One line of methods strengthens the retrieval pipeline while keeping the context size essentially fixed. Classical sparse retrievers and dense dual-encoder retrievers improve the recall and coarse ranking of candidate passages, and are often coupled with cross-encoder or sequence-to-sequence rerankers that refine the fine-grained ordering of documents (robertson2009probabilistic; karpukhin-etal-2020-dense; xiong2020approximatenearestneighbornegative; nogueira2020documentrankingpretrainedsequencetosequence). More recently, LLM-based rerankers score and prune contexts using query-aware, list-aware, or generator-aware signals (sun2024chatgptgoodsearchinvestigating; chen2025scirerankbenchbenchmarkingrerankersscientific; drozdov2023paradepassagerankingusing; wang2024learningretrieveincontextexamples; deng2025influenceguidedcontextselection). These approaches are effective at promoting highly relevant passages in the ranked list and demoting obvious distractors, but the generator still typically consumes either a fixed top-$K$ prefix or a set obtained by hand-crafted thresholds, so the fundamental trade-off between missing evidence and accumulating noise remains.
A complementary line of work explicitly adapts the number of retrieved passages. Adaptive-RAG routes each query to no-retrieval, single-step, or iterative RAG pipelines based on a learned complexity classifier (jeong2024adaptiveraglearningadaptretrievalaugmented), while adaptive-$k$ methods select the cutoff $K$ from the similarity-score distribution of the retrieved candidates without additional model tuning or extra LLM calls (taguchi2025efficientcontextselectionlongcontext). Although such methods alleviate the mismatch between simple and complex queries, they still rely on heuristic decision rules over per-passage similarity scores, and do not directly optimize for a minimal sufficient evidence subset under a given token budget.
@@
To move beyond fixed heuristics, reinforcement learning (RL) has recently been explored as a way to optimize retrieval and selection policies directly from task feedback while keeping test-time inference to a single policy forward pass. DynamicRAG models the reranker as an RL agent over document sequences and uses LLM-judged answer quality as reward to jointly adjust both the order and the number of retrieved documents (sun2025dynamicragleveragingoutputslarge). Beyond reranking, recent RL-based systems such as Memory-R1 and related memory agents frame long-term memory management and retrieval decisions as RL problems, training policies to decide what to store, update, or retrieve in order to support downstream QA and dialogue (yan2025memoryr1enhancinglargelanguage). RL has also been applied to conversational query reformulation and retrieval alignment and to broader agentic RAG frameworks that optimize multi-step retrieval and reasoning trajectories (zhu2025convsearchr1enhancingqueryreformulation; xiong2025raggymsystematicoptimizationlanguage; jiang2025rexragreasoningexplorationpolicy). However, existing RL-style approaches still suffer from largely trajectory-level and sparse rewards, which makes it difficult to assign credit to individual passages or penalize redundancy, and they are typically trained to improve list-wise ranking quality or memory operations rather than to identify a minimal evidence subset that preserves answerability under a fixed input budget.
@@
To address these challenges, we introduce Context-Picker, a reasoning-aware framework that fundamentally shifts the context selection paradigm from similarity-based ranking to minimal sufficient subset selection. Instead of treating retrieval as a sorting problem, Context-Picker formulates it as a decision-making process, learning to construct a variable-length evidence set that is strictly necessary for answering the query. Central to our approach is a human-inspired Coarse-to-Fine optimization strategy implemented via a two-stage reinforcement learning schedule. In Stage I (Recall-Oriented), the picker is trained to maximize information with a relaxed redundancy margin, so that all potentially relevant reasoning chains—especially those spanning multiple passages—are captured. In Stage II (Precision-Oriented), the objective shifts to refinement: the policy learns to prune redundant or weakly relevant passages, distilling the context into a compact, noise-free subset without compromising answerability. To stabilize training and alleviate reward sparsity, we introduce an offline evidence distillation pipeline that uses a generator–judge loop with greedy Leave-One-Out (LOO) pruning to mine “minimal sufficient” evidence sets from raw documents. These distilled sets provide dense, task-aligned supervision, enabling the policy to learn the contribution of each evidence piece. Extensive experiments on five long-context and multi-hop QA benchmarks demonstrate that Context-Picker significantly outperforms strong RAG baselines, achieving superior answer accuracy with comparable or reduced context lengths.
@@
Our main contributions are summarized as follows:
  • We propose Context-Picker, a reasoning-aware context picker trained with a two-stage reinforcement learning scheme and redundancy-aware reward shaping. The picker jointly decides which passages to keep and how many to include, with a recall-oriented stage for high-coverage picking and a precision-oriented stage for aggressive compression, explicitly addressing the limitations of fixed top-$K$ selection in long-context QA.
  • We introduce an offline evidence mining pipeline that mines greedily minimal sufficient evidence sets via a generator–judge loop and a leave-one-out pruning procedure, providing high-quality, task-aligned supervision for training the picker.
  • We conduct extensive experiments on five long-context and multi-hop QA benchmarks, showing that Context-Picker improves LLM-as-judge accuracy over strong RAG baselines on four datasets and achieves favorable accuracy–efficiency trade-offs on the remaining one, with ablations validating the impact of each key component.
@@ 2.1 Retrieval-Augmented Generation
We follow the standard retrieval-augmented generation (RAG) formulation (lewis2021retrievalaugmentedgenerationknowledgeintensivenlp; guu2020realmretrievalaugmentedlanguagemodel; izacard2022atlasfewshotlearningretrieval). Let $\mathcal{D}$ denote a large non-parametric corpus (e.g., Wikipedia or a long-term memory store). In long-context QA, each document in $\mathcal{D}$ is first segmented into shorter passages (“chunks”), which serve as the retrieval units. Given a query $q$, a retriever operates over these passages and returns a candidate pool of at most $K_{\max}$ passages
@@
optionally refined by a reranker that reorders $\mathcal{C}(q)$ according to query-specific relevance (karpukhin-etal-2020-dense; xiong2020approximatenearestneighbornegative; nogueira2020documentrankingpretrainedsequencetosequence). Unless otherwise stated, we use $\mathcal{C}(q)$, or simply $\mathcal{C}$ when the query is clear from context, to denote this (re)ranked candidate pool. Later, when formulating Context-Picker, we additionally attach a unique identifier to each passage $c_{j}$ and write $\mathcal{C}=\{(c_{j},\text{id}_{j})\}_{j=1}^{N}$ for convenience.
@@
$\mathcal{S}^{\star}\;\in\;\arg\max_{\mathcal{S}\subseteq\mathcal{C}(q)}\Bigl(U(q,\mathcal{S})\;-\;\lambda\cdot\mathrm{Len}(\mathcal{S})\Bigr)\quad\text{s.t.}\quad\mathrm{Tok}(q,\mathcal{S})\leq B,$ (2)
where $U(q,\mathcal{S})$ is a task utility, $\mathrm{Len}(\mathcal{S})$ measures the size of the support set, $\lambda\geq 0$ controls the quality–brevity trade-off, and $\mathrm{Tok}(q,\mathcal{S})$ counts input tokens. A common baseline uses a fixed top-$K$ prefix $\mathcal{S}=\{\tilde{c}_{1},\ldots,\tilde{c}_{K}\}$ for all queries, which under- or over-includes context depending on query difficulty and can suffer from “lost in the middle” effects in long prompts (liu2023lostmiddlelanguagemodels; jin2024longcontextllmsmeetrag). Adaptive strategies instead learn a policy $\pi_{\phi}(\mathcal{S}\mid q,\mathcal{C}(q))$ that jointly decides which passages to keep and how many to include (sun2025dynamicragleveragingoutputslarge; deng2025influenceguidedcontextselection). Context-Picker builds on this formulation and learns a reasoning-aware policy under a token budget.
@@
Figure 2: Overview of the Context-Picker framework. The pipeline consists of two parts: (1) Offline Evidence Mining, where a generator–judge loop employs a Leave-One-Out (LOO) strategy to mine minimal sufficient evidence sets ($\mathcal{S}_{\text{gold}}$) as supervision; and (2) Context-Picker Pipeline, where the picker policy ($\pi_{\theta}$) learns to select evidence from retrieved candidates ($\mathcal{C}$). The training follows a Coarse-to-Fine schedule: Stage I optimizes for high recall to capture reasoning chains, while Stage II tightens the redundancy penalty to distill a compact support set, guided by GRPO updates.
2.2 Group Relative Policy Optimization (GRPO)
We view evidence picking as a policy optimization problem. Let $o$ denote an observation which consists of a query $q$ and its candidate pool $\mathcal{C}(q)$, and let $a$ denote a discrete action (a set of picked passage IDs). A stochastic policy $\pi_{\phi}(a\mid o)$ with parameters $\phi$ induces the objective
$J(\phi)\;=\;\mathbb{E}_{o\sim\mathcal{D}_{\text{train}},\,a\sim\pi_{\phi}(\cdot\mid o)}\bigl[\,R(a,o)\,\bigr],$ (3)
where $R(a,o)$ is a task-specific reward.
We used Group Relative Policy Optimization (GRPO) (shao2024deepseekmathpushinglimitsmathematical) to optimize our training goal. For each observation $o$ (e.g., a query and its candidate pool), the policy $\pi_{\phi}$ (with a frozen reference policy $\pi_{\phi_{\mathrm{old}}}$) samples a group of $G$ candidate actions $\{a_{i}\}_{i=1}^{G}\sim\pi_{\phi_{\mathrm{old}}}(\cdot\mid o),$ and each action receives a scalar reward $R_{i}=R(a_{i},o).$
The group-normalized advantage for the $i$-th action is $\hat{A}_{i}=\frac{R_{i}-\operatorname{mean}\bigl(\{R_{j}\}_{j=1}^{G}\bigr)}{\operatorname{std}\bigl(\{R_{j}\}_{j=1}^{G}\bigr)+\epsilon},$ where $\epsilon$ is a small constant for numerical stability.
The probability ratio is defined as
$r_{i}(\phi)=\frac{\pi_{\phi}(a_{i}\mid o)}{\pi_{\phi_{\mathrm{old}}}(a_{i}\mid o)}.$ (4)
Our GRPO objective with decoupled, asymmetric clipping is
$\begin{split}J_{\mathrm{GRPO}}(\phi)&=\mathbb{E}_{o\sim\mathcal{D},\;\{a_{i}\}_{i=1}^{G}\sim\pi_{\phi_{\mathrm{old}}}(\cdot\mid o)}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\min\Big(r_{i}(\phi)\,\hat{A}_{i},\\ &\qquad\qquad\operatorname{clip}\big(r_{i}(\phi),\,1-\epsilon_{\mathrm{low}},\,1+\epsilon_{\mathrm{high}}\big)\,\hat{A}_{i}\Big)\Bigg]-\beta\cdot\mathrm{KL}\!\left(\pi_{\phi}\,\|\,\pi_{\phi_{\mathrm{old}}}\right),\end{split}$ (5)
where $\epsilon_{\mathrm{low}},\epsilon_{\mathrm{high}}>0$ control the asymmetric clipping range and $\beta\geq 0$ controls the KL regularization strength.
@@ 3 Context-Picker
In this section, we present Context-Picker, our reinforcement learning–based context picker. An overview of the framework is shown in Figure 2. We first formulate context picking as a single-step Markov decision process (MDP) in Section 3.1. We then describe the overall framework, which consists of two components: (i) an offline evidence mining pipeline (Section 3.2) that distills minimal sufficient evidence sets from raw documents, and (ii) a multi-stage reinforcement learning procedure (Section 3.3) that trains a picker policy with a recall-oriented stage followed by a precision-oriented stage. Finally, we detail how the learned picker is integrated with the downstream generator at inference time, and summarize the resulting inference pipeline in Algorithm 3.3.

### v2

@@ Abstract
In long-context question answering, selecting the appropriate scope of context for a query remains a key and unresolved challenge. Insufficient context can lead to missing essential information, whereas excessive context often introduces noise and degrades answer quality. Conventional methods, such as retrieving a fixed number of passages or applying reranking, struggle to dynamically determine which context to include. This is especially problematic for factoid questions, which typically depend only on a few precise pieces of evidence. To overcome this limitation, we propose Context-Picker, a reasoning-aware framework that reframes context selection as the task of identifying a minimal sufficient evidence subset, moving beyond conventional similarity-based ranking. Context-Picker uses a human-inspired two-stage reinforcement learning schedule: stage 1 focuses on improving the recall rate of critical passages, and stage 2 prioritizes pruning redundancy to distill a compact evidence set. To resolve reward sparsity, we propose an offline evidence distillation pipeline that mines “minimal sufficient sets” via a Leave-One-Out (LOO) procedure, providing dense and task-aligned supervision. Experiments on five long-context and multi-hop QA datasets demonstrate that our method outperforms strong RAG baselines and achieved higher answer accuracy. Ablation studies also indicate that our coarse-to-fine optimization schedule, the redundancy-aware reward shaping, along with the rationale generated by the policy, all contribute substantially to these gains.
@@ 1 Introduction
In long-context question answering (LCQA) tasks, Retrieval-Augmented Generation (RAG) has become a standard paradigm for Large Language Models (LLMs) to retrieve knowledge beyond their parametric knowledge Lewis et al. (2021); Guu et al. (2020); Izacard et al. (2022). RAG enables LLMs to access query-related information and mitigates hallucination. In practice, most RAG systems adopt a fixed Top-$K$ strategy: a retriever ranks passages based on their semantic similarity with the query, and only the top-$K$ passages with highest similarity scores are concatenated and fed to the LLMs. Although this method is generally effective, it still faces the problem of identifying the suitable $K$ for each query. If $K$ is too small, the model may miss critical evidence, leading to failure in answering the question. If $K$ is overly large, too many weakly related passages will result in degrading answer quality through distractors, attention dilution, and the ”lost-in-the-middle” phenomenon, where LLMs underutilize information placed in the middle of long prompts Liu et al. (2023). As shown in Figure 1, increasing retrieval depth monotonically improves recall but often yields diminishing returns for answer accuracy Jin et al. (2024), which suggests that context handling should be framed not as ranking, but as subset selection–identifying a compact, query-specific evidence set sufficient for reasoning.
@@
Figure 1: Accuracy versus retrieval depth (Top-$K$) in a standard RAG pipeline. Recall increases with $K$, but answer accuracy does not improve, which is also reported in recent RAG studies Jin et al. (2024).
@@
Recent studies on retrieval-augmented generation address this problem from two principal approaches. The first line of work enhances the retrieval pipeline while largely maintaining a fixed context size. Conventional sparse retrievers and dense dual-encoder models improve the recall using coarse ranking of candidate passages, and are frequently combined with cross-encoder or sequence-to-sequence rerankers that further refine the fine-grained ordering of documents Robertson et al. (2009); Karpukhin et al. (2020); Xiong et al. (2020); Nogueira et al. (2020). More recent efforts employ LLM-based rerankers, which leverage query-aware, list-aware, or generator-aware signals to score and filter contexts Sun et al. (2024); Chen et al. (2025); Drozdov et al. (2023); Wang et al. (2024); Deng et al. (2025). While these methods are effective at elevating highly relevant passages and suppressing obvious distractors in the ranked list, the generator typically still processes either a fixed top-K set or manually defined thresholds; thus, the inherent trade-off between missing pertinent evidence and accumulating irrelevant noise still exists.
A complementary line of work adapts how much external evidence to use on a per-query basis. Adaptive-RAG trains a lightweight classifier to route each query to different retrieval depths (e.g., no retrieval, single-step retrieval, or iterative retrieval) based on estimated question complexity Jeong et al. (2024). In contrast, Adaptive-$k$ selects a query- and context-specific cutoff $K$ by thresholding the similarity-score distribution of retrieved candidates, requiring neither additional fine-tuning nor extra LLM calls Taguchi et al. (2025). While these approaches avoid using a fixed $k$, they rely on either a learned query-level router or on non-learned score-thresholding rules, and they do not explicitly optimize for identifying a non-redundant evidence subset that preserves answerability under a fixed input budget.
@@
To move beyond routing and thresholding rules, reinforcement learning (RL) has recently been explored as a way to optimize retrieval and selection policies directly from task feedback while keeping test-time inference to a single policy forward pass. DynamicRAG models the reranker as an RL agent over document sequences and uses LLM-judged answer quality as reward to jointly adjust both the order and the number of retrieved documents Sun et al. (2025). Beyond reranking, recent RL-based systems such as Memory-R1 and related memory agents frame long-term memory management and retrieval decisions as RL problems, where a policy is trained to decide what to store, update, or retrieve in order to support downstream QA and dialogue Yan et al. (2025). RL has also been applied to conversational query reformulation, retrieval alignment and agentic RAG frameworks that optimize multi-step retrieval and reasoning trajectories Zhu et al. (2025); Xiong et al. (2025); Jiang et al. (2025). However, many existing RL-style approaches still rely on trajectory-level and relatively sparse rewards, making it difficult to credit assignment to individual passages and suppress redundant content. Moreover, they are often optimized for ranking quality, memory operations, or trajectory success rather than for explicitly finding the smallest evidence set that maintains answerability under a strict token budget.
@@
In this paper, we introduce Context-Picker, a reasoning-aware framework that fundamentally shifts the context selection paradigm from similarity-based ranking to minimal sufficient subset selection. Instead of treating retrieval as a sorting problem, Context-Picker formulates it as a decision-making process, and learns to construct a variable-length evidence set that is strictly necessary for answering the query. Central to our approach is a human-inspired Coarse-to-Fine optimization strategy implemented via a two-stage reinforcement learning schedule. In Stage I (Recall-Oriented), the picker is trained to maximize information with a relaxed redundancy margin, so that all potentially relevant passages—especially those spanning multiple passages—are captured. In Stage II (Precision-Oriented), the objective shifts to refinement: the policy learns to prune redundant or weakly relevant passages, distilling the context into a compact, noise-free subset without compromising answerability. To alleviate reward sparsity and stabilize training, we introduce an offline evidence distillation pipeline that uses a generator-judge loop with greedy Leave-One-Out (LOO) pruning to mine minimal sufficient evidence sets from raw documents. These distilled sets provide dense, task-aligned supervision, enabling the picker (i.e., policy) to learn to measure the contribution of each evidence piece. Experiments on five long-context and multi-hop QA datasets demonstrate that our method outperforms strong RAG baselines and achieved higher answer accuracy.
@@
Our main contributions are three-fold: (1) We propose Context-Picker, a reasoning-aware context picker trained with a two-stage reinforcement learning scheme (Recall-Oriented and Precision-Oriented) to jointly learn which and how many passages to pick. (2) We introduce an offline evidence mining pipeline that uses a generator-judge loop with Leave-One-Out (LOO) pruning to create minimal sufficient evidence sets for dense supervision. (3) We conduct extensive experiments on five long-context and multi-hop QA datasets, demonstrating that Context-Picker improves accuracy over strong RAG baselines on four out of five datasets.
Figure 2: Overview of the Context-Picker framework. The pipeline consists of two parts: (1) Offline Evidence Mining, where a generator–judge loop employs a Leave-One-Out (LOO) strategy to mine minimal sufficient evidence sets ($\mathcal{S}_{\text{gold}}$) as supervision; (2) Context-Picker Pipeline, where the picker policy ($\pi_{\theta}$) learns to select evidence from retrieved candidates ($\mathcal{C}$). The training follows a Coarse-to-Fine schedule: Stage I optimizes for high recall to capture reasoning chains, while Stage II tightens the redundancy penalty to distill a compact support set, guided by GRPO updates.
@@ 2.1 Retrieval-Augmented Generation
We follow the standard retrieval-augmented generation (RAG) formulation Lewis et al. (2021); Guu et al. (2020); Izacard et al. (2022). Let $\mathcal{D}$ denote a large non-parametric corpus (e.g., Wikipedia or a long-term memory store). In long-context QA, each document in $\mathcal{D}$ is first segmented into shorter passages (“chunks”), which serve as the retrieval units. Given a query $q$, the retriever returns the top-$K$ most relevant passages.
@@
optionally refined by a reranker that reorders $\mathcal{C}(q)$ according to query-specific relevance Karpukhin et al. (2020); Xiong et al. (2020); Nogueira et al. (2020). We use $\mathcal{C}(q)$, or simply $\mathcal{C}$ when the query is clear from context, to denote this (re)ranked candidate pool. Later, when formulating Context-Picker, we additionally attach a unique identifier to each passage $c_{j}$ and write $\mathcal{C}=\{(c_{j},\text{id}_{j})\}_{j=1}^{N}$ for convenience.
@@
$\mathcal{S}^{\star}$ $\in\arg\max_{\mathcal{S}\subseteq\mathcal{C}(q)}\Bigl(U(q,\mathcal{S})-\lambda\cdot\mathrm{Len}(\mathcal{S})\Bigr)$ (2)
$\quad\text{s.t.}\quad\mathrm{Tok}(q,\mathcal{S})\leq B.$
where $U(q,\mathcal{S})$ is a task utility, $\mathrm{Len}(\mathcal{S})$ measures the size of the support set, $\lambda\geq 0$ controls the quality–brevity trade-off, and $\mathrm{Tok}(q,\mathcal{S})$ counts input tokens. Context-Picker builds on this formulation and learns a reasoning-aware policy under a token budget.
@@
2.2 Optimization Objective
We employ Group Relative Policy Optimization (GRPO) Shao et al. (2024) to train the picker. Unlike PPO, GRPO eliminates the critic model by estimating baselines from group averages. For each query $q$, GRPO samples a group of outputs $\{a_{i}\}_{i=1}^{G}$ from the old policy $\pi_{\theta_{\mathrm{old}}}$:
$\tiny J(\theta)=\mathbb{E}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(r_{i}(\theta)\hat{A}_{i},\text{clip}(r_{i}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i}\right)\right]-\beta D_{KL},$ (3)
where $r_{i}(\theta)=\pi_{\theta}(a_{i}|q)/\pi_{\theta_{\mathrm{old}}}(a_{i}|q)$ and the advantage $\hat{A}_{i}$ is the normalized reward within the group.
@@ 3 Context-Picker
In this section, we introduce Context-Picker, a reinforcement learning based method for selecting context passages. Figure 2 provides an overview of the proposed framework. We start by formulating context picking as a single-step Markov decision process (MDP) in Section 3.1. Building on this formulation, we detail our overall framework, which comprises two key components: (i) an offline evidence mining pipeline (Section 3.2) that extracts compact yet sufficient evidence sets from raw documents, and (ii) a multi-stage reinforcement learning scheme (Section 3.3) that learns a picker policy via a recall-oriented stage followed by a precision-oriented stage. Finally, we detail how the learned picker is integrated with the downstream generator at inference time.